### v1

@@ 1 Introduction
Kaczmarz(-type) methods have proven effective when the linear system is highly over-determined and the computing environment restricts access to the input data, thus benefiting from their cheap and localized updates. However, the popular belief is that, outside of these considerations, Krylov subspace methods such as Conjugate Gradient (CG) [HS52], LSQR [PS82], and GMRES [SS86] are typically superior to Kaczmarz methods on account of their ability to exploit outliers in the input’s singular value distribution, attaining fast convergence even for some highly ill-conditioned systems [AL86]. In this work, we re-examine this assertion, developing Kaczmarz methods that can similarly exploit outlying singular values in the input to achieve fast convergence, going even beyond what Krylov subspace methods can attain for a natural class of singular value distributions. Crucially, our proposed methods do not require the systems to be very tall or even over-determined to work well.
@@
where $\tilde{O}$ hides logarithmic factors described in detail alongside the theorem, while $\bar{\kappa}_{k}$ is the normalized Demmel condition number of the matrix ${\mathbf{A}}$ excluding the top-$k$ part of its singular value decomposition (SVD):
@@
Here, ${\mathbf{A}}_{k}=\sum_{i=1}^{k}\sigma_{i}{\mathbf{u}}_{i}{\mathbf{v}}_{i}^{\scriptscriptstyle{\top}}$ denotes the top-$k$ part of ${\mathbf{A}}$’s SVD, and $\|\cdot\|_{F}$ is the matrix Frobenius norm. Demmel condition number $\|{\mathbf{M}}\|_{F}\|{\mathbf{M}}^{\dagger}\|=\|{\mathbf{M}}\|_{F}/\sigma_{\min}^{+}({\mathbf{M}})$ is often used to describe the convergence rate of Kaczmarz methods, starting from [SV09], and when normalized by $\sqrt{\mathrm{rank}({\mathbf{M}})}$, it is always upper-bounded by the classical condition number $\kappa({\mathbf{M}}):=\|{\mathbf{M}}\|\|{\mathbf{M}}^{\dagger}\|$.
As suggested by (1.1), Kaczmarz++ exhibits two phases of convergence: In the first phase, the algorithm implicitly captures the top-$k$ part of ${\mathbf{A}}$’s spectrum via our block memoization scheme that requires only $\tilde{O}(nk)$ additional memory. Then, in the second phase, it uses this information (along with our adaptive acceleration scheme) to attain a convergence rate that is independent of the top-$k$ singular values of ${\mathbf{A}}$.
To put this result in context, consider a comparable convergence guarantee achievable by a Krylov subspace method (such as LSQR) for solving a dense $n\times n$ ill-conditioned linear system with $k$ large outlying singular values. Such a method is also expected to exhibit two phases of convergence, where the first phase builds the Krylov subspace that captures the top-$k$ part of ${\mathbf{A}}$’s SVD, while the second phase leverages it to attain fast convergence, reaching $\epsilon$ error after:
@@
Here, $\kappa_{k}\!:=\!\kappa({\mathbf{A}}-{\mathbf{A}}_{k})$ is the condition number of ${\mathbf{A}}$ excluding its top-$k$ singular values.
There are two key differences between the guarantees (1.1) and (1.3). First, the fast convergence of Kaczmarz++ relies on $\bar{\kappa}_{k}$, which can be substantially smaller than $\kappa_{k}$. Second, and more importantly, the first phase of Kaczmarz++ takes only $\tilde{O}(n^{2}+nk^{2})$ time, which is much more efficient than the $O(n^{2}k)$ first phase of Krylov when $k$ is sufficiently larger than a constant (and smaller than $n$). This is because building the Krylov subspace for the outlying singular values requires at least $k$ matrix-vector products with the full matrix ${\mathbf{A}}$, whereas in the initial phase of Kaczmarz++, the algorithm iterates over $\tilde{O}(n/k)$ blocks of $\tilde{O}(k)$ rows each, a computational equivalent of only a few matrix-vector products.
Naturally, the above guarantees do not provide a complete convergence comparison between Kaczmarz++ and Krylov subspace methods (e.g., they do not account for small outlying singular values). However, they indicate that Kaczmarz methods can be competitive with Krylov solvers in terms of arithmetic operations even for square linear systems, and not only for highly over-determined ones, as is often suggested. To verify this claim empirically, we develop a practical implementation of Kaczmarz++ that is specifically optimized for positive semidefinite systems (CD++, Algorithm 3), and test it on a collection of benchmark problems from the machine learning literature [VvRBT13, PVG+11], which are known to exhibit large outlying eigenvalues. The experiments confirm our theoretical findings, showing that our algorithm is competitive in floating point operations with both CG and GMRES on a range of input matrices.
@@ 1.1 Overview of the Main Algorithm
Consider solving a consistent linear system with $m$ equations, ${\mathbf{A}}{\mathbf{x}}={\mathbf{b}}$, where ${\mathbf{A}}\in{\mathds{R}}^{m\times n}$ and ${\mathbf{b}}\in{\mathds{R}}^{n}$. The classical block Kaczmarz method constructs a sequence of iterates ${\mathbf{x}}_{0},{\mathbf{x}}_{1},...$ by repeatedly choosing a subset $S=S(t)\subseteq\{1,...,m\}=:[m]$ of those equations, and then projecting the current iterate ${\mathbf{x}}_{t}$ onto the subspace of solutions of those equations, i.e., the under-determined system ${\mathbf{A}}_{S}{\mathbf{x}}={\mathbf{b}}_{S}$. This leads to the block Kaczmarz update which can be stated as follows:
@@
We note that the Hadamard matrix, similarly to a Discrete Fourier Transform (DFT), is a scaled orthogonal matrix (specifically, ${\mathbf{H}}^{\scriptscriptstyle{\top}}{\mathbf{H}}=m{\mathbf{I}}_{m}$) that admits fast matrix-vector multiply (see Appendix C for a detailed discussion). In fact, these are the only two properties we use in our analysis, and one could replace Hadamard with a DFT or other fast transforms.
Since ${\mathbf{Q}}^{\scriptscriptstyle{\top}}{\mathbf{Q}}={\mathbf{I}}$, transforming the system ${\mathbf{A}}{\mathbf{x}}={\mathbf{b}}$ into ${\mathbf{Q}}{\mathbf{A}}{\mathbf{x}}={\mathbf{Q}}{\mathbf{b}}$ does not affect the solution of the system nor does it affect the singular value distribution of the input matrix. However, crucially, it ensures that all equations become roughly equally important (this is known as incoherence), which allows us to select a representative subset $S$ uniformly at random. Using the fast matrix-vector multiply, this transformation can be done using $mn\log m$ arithmetic operations. This preprocessing step, while essential for parts of our analysis, may be skipped when the input matrix naturally exhibits an incoherence property [NT14].
@@
1:Input: ${\mathbf{A}}\in{\mathds{R}}^{m\times n}$, ${\mathbf{b}}\in{\mathds{R}}^{n}$, block size $s$, iterate ${\mathbf{x}}_{0}$, parameters $B$, ${\lambda}$, $\rho$, $\eta$;
@@ 1.2 Related work
Recently, there has been a number of works suggesting that variants of block Kaczmarz can exploit $k$ large outliers in the singular value distribution, by expressing its convergence in terms of $\bar{\kappa}_{k}$ or $\kappa_{k}$. [DR24] were the first to show this, however due to the large cost of Gaussian sketching and lack of acceleration, their method does not have a computational benefit over existing approaches. Then, [DY24] showed that a similar convergence guarantee can be obtained by block Kaczmarz with uniformly sampled blocks, after RHT preprocessing. Their algorithm converges in $\tilde{O}((n^{2}+nk^{2})\bar{\kappa}_{k}^{2}\log 1/\epsilon)$ operations. Most recently, [DLNR24] obtained $\tilde{O}((n^{2}+nk^{2})\kappa_{k}\log 1/\epsilon)$ operations by introducing momentum acceleration via the scheme of [GHRS18], but at the expense of relying on sparse sketching for block construction (much slower than uniform sampling).
@@ 1.3 Notation
We let $[m]:=\{1,...,m\}$, whereas (${[m]\choose s}$ denotes all size $s$ subsets of $[m]$. For a matrix ${\mathbf{A}}\in{\mathds{R}}^{m\times n}$ and subset $S\in{[m]\choose s}$, we use ${\mathbf{A}}_{S}\in{\mathds{R}}^{s\times n}$ to denote the submatrix of the rows of ${\mathbf{A}}$ indexed by $S$, and if $m=n$, then ${\mathbf{A}}_{S,S}\in{\mathds{R}}^{s\times s}$ is the principal submatrix indexed by $S$. We use $\|{\mathbf{A}}\|$, $\|{\mathbf{A}}\|_{F}$, and ${\mathbf{A}}^{\dagger}$ to denote the spectral/Frobenius norms of matrix ${\mathbf{A}}$ as well as its Moore-Penrose pseudoinverse, while $\lambda_{\max}({\mathbf{A}})$ and $\lambda_{\min}^{+}({\mathbf{A}})$ are the largest and smallest positive eigenvalues of an $n\times n$ positive semidefinite (PSD) matrix, denoted ${\mathbf{A}}\in\mathcal{S}_{n}^{+}$. For ${\mathbf{A}}\in\mathcal{S}_{n}^{+}$ and ${\mathbf{v}}\in{\mathds{R}}^{n}$, we use $\|{\mathbf{v}}\|_{{\mathbf{A}}}=\sqrt{{\mathbf{v}}^{\scriptscriptstyle{\top}}{\mathbf{A}}{\mathbf{v}}}$, and for symmetric matrices, ${\mathbf{A}}\preceq{\mathbf{B}}$ means that ${\mathbf{B}}-{\mathbf{A}}\in\mathcal{S}_{n}^{+}$. We use $C>0$ to denote an absolute constant, which may change from line to line.
@@ 2 Stable Convergence with Adaptive Acceleration
Given ${\mathbf{A}}\in{\mathds{R}}^{m\times n}$ and ${\mathbf{b}}\in{\mathds{R}}^{n}$, let ${\mathbf{x}}^{*}$ be the minimum-norm solution to ${\mathbf{A}}{\mathbf{x}}={\mathbf{b}}$. Also, let $\mathcal{D}$ be a probability distribution over subsets of $[m]$. Given $\lambda\geq 0$, define the random regularized projection:
@@
Let $\rho\in[0,c\bar{\rho}]$ and $\eta\in[\frac{c}{\nu},\frac{1}{2\nu}]$ for some $c\in(0,1/2]$, and suppose that a sequence ${\mathbf{x}}_{t}$ is updated as in lines 6-10 of Algorithm 1, allowing the regularized projection step (line 8) to be computed inexactly, returning $\tilde{\mathbf{w}}_{t}$ such that $\|\tilde{\mathbf{w}}_{t}-{\mathbf{w}}_{t}\|\leq\frac{\rho^{2}}{8\eta}\|{\mathbf{w}}_{t}\|$. Then,
@@
If $\|\tilde{\mathbf{w}}_{t}-{\mathbf{w}}_{t}\|\leq\frac{\tilde{\mu}}{4}\|{\mathbf{w}}_{t}\|$, then we have
@@ 3.1 Expectation of the Regularized Projection
Suppose ${\mathbf{A}}\in{\mathds{R}}^{m\times n}$ is transformed by RHT, i.e., $\bar{\mathbf{A}}={\mathbf{Q}}{\mathbf{A}}$. Let $\sigma_{1}\geq\sigma_{2}\geq...$ be ${\mathbf{A}}$’s singular values. Given $\delta\in(0,1)$ and $C\log(m/\delta)\leq k<\mathrm{rank}({\mathbf{A}})$, let $\bar{\lambda}=\frac{1}{k}\sum_{i>k}\sigma_{i}^{2}$. Let $S\sim{\mathcal{U}}(m,s)$ be a uniformly random subset of $[m]$ with size $s\geq Ck\log(m\bar{\kappa}_{k})$. Then, for any $0\leq{\lambda}\leq\frac{k}{m}\bar{\lambda}$, with probability $1-\delta$ the transformed matrix $\bar{\mathbf{A}}$ satisfies:
@@
$\mu\Big{(}\bar{\mathbf{A}},{\mathcal{U}}(m,s),\lambda\Big{)}\geq\frac{\sigma_{\min}^{+}({\mathbf{A}})^{2}/2}{\sigma_{\min}^{+}({\mathbf{A}})^{2}+\bar{\lambda}}=\frac{1/2}{1+\frac{r-k}{k}\bar{\kappa}_{k}^{2}}\geq\frac{k}{2r\bar{\kappa}_{k}^{2}}\geq\frac{k}{2m\bar{\kappa}_{k}^{2}},$
@@
Conditioned on the event $\mathcal{A}_{0}:=[S_{\mathrm{DPP}}\subseteq S]$ (which holds with probability $1-\delta^{\prime}$), we have ${\mathbf{A}}_{S_{\mathrm{DPP}}}^{\scriptscriptstyle{\top}}{\mathbf{A}}_{S_{\mathrm{DPP}}}\preceq{\mathbf{A}}_{S}^{\scriptscriptstyle{\top}}{\mathbf{A}}_{S}$. Combining this with (3.1) and Lemma 3.6, we have the following holds:
${\mathds{E}}[{\mathbf{P}}_{{\lambda},S}]=$ $~{}{\mathds{E}}[{\mathbf{P}}_{{\lambda},S}\mid\mathcal{A}_{0}]\Pr\{\mathcal{A}_{0}\}+{\mathds{E}}[{\mathbf{P}}_{{\lambda},S}\mid\mathcal{A}_{0}^{c}]\Pr\{\mathcal{A}_{0}^{c}\}$
$\succeq$ $~{}{\mathds{E}}[{\mathbf{P}}_{{\lambda},S}\mid\mathcal{A}_{0}]\Pr\{\mathcal{A}_{0}\}$
$\succeq$ $~{}\Pr\{\mathcal{A}_{0}\}\cdot{\mathbf{I}}-\frac{k\bar{\lambda}}{m}\cdot{\mathds{E}}\left[\left({\mathbf{A}}_{S}^{\scriptscriptstyle{\top}}{\mathbf{A}}_{S}+\frac{k\bar{\lambda}}{m}{\mathbf{I}}\right)^{-1}\biggl{|}\mathcal{A}_{0}\right]\cdot\Pr\{\mathcal{A}_{0}\}$
$\succeq$ $~{}\Pr\{\mathcal{A}_{0}\}\cdot{\mathbf{I}}-{\mathds{E}}\left[\left(\frac{m}{k\bar{\lambda}}{\mathbf{A}}_{S_{\mathrm{DPP}}}^{\scriptscriptstyle{\top}}{\mathbf{A}}_{S_{\mathrm{DPP}}}+{\mathbf{I}}\right)^{-1}\biggl{|}\mathcal{A}_{0}\right]\cdot\Pr\{\mathcal{A}_{0}\}$

### v2

@@ 1 Introduction
Kaczmarz(-type) methods have proven effective when the linear system is highly over-determined and the computing environment restricts access to the input data, thus benefiting from their cheap and localized updates. However, outside of these considerations, Krylov subspace methods such as Conjugate Gradient (CG) [HS52], LSQR [PS82], and GMRES [SS86] typically appear (theoretically) superior to Kaczmarz methods on account of their ability to exploit outliers and clusters in the input’s singular value distribution, attaining fast convergence even for some highly ill-conditioned systems (e.g., see Chapter 5 of [NW99]). In this work, we re-examine this assertion, developing Kaczmarz methods that can similarly exploit outlying singular values in the input to achieve fast convergence, going even beyond what Krylov subspace methods can attain for a natural class of singular value distributions. Crucially, our proposed methods do not require the systems to be very tall or even over-determined to work well.
@@
[…]
@@
[…]
There are two key differences between the guarantees (1.1) and (1.3). First, the fast convergence of Kaczmarz++ relies on $\bar{\kappa}_{k}$, which can be substantially smaller than $\kappa_{k}$. Second, and more importantly, the first phase of Kaczmarz++ takes only $\tilde{O}(n^{2}+nk^{2})$ time, which can be much more efficient than the $O(n^{2}k)$ first phase of Krylov. This is because building the Krylov subspace for the outlying singular values requires at least $k$ matrix-vector products with the full matrix ${\mathbf{A}}$. Lower bounds show that this cost is necessary for any algorithm based solely on matrix-vector products [DLNR24] (including all Krylov methods) when solving systems with $k$ large outlying singular values. On the other hand, in its initial phase, Kaczmarz++ leverages direct access to ${\mathbf{A}}$ by iterating over $\tilde{O}(n/k)$ blocks of $\tilde{O}(k)$ rows each, a computational equivalent of only a few matrix-vector products. This benefit of Kaczmarz++ is significant when $k$ is sufficiently larger than a logarithmic power of $n$ and sufficiently smaller than $n$.
Naturally, the above guarantees do not provide a complete convergence comparison between Kaczmarz++ and Krylov subspace methods (e.g., they do not account for small outlying singular values). However, they indicate that Kaczmarz methods can work well on ill-conditioned problems and are competitive with Krylov solvers in terms of arithmetic operations even for square linear systems, and not only for highly over-determined ones, as is often suggested. To verify these claims empirically, we develop a practical implementation of Kaczmarz++, performing an ablation study on ill-conditioned synthetic matrices with large outlying singular values. We also develop another variant of our algorithm which is specifically optimized for positive semidefinite systems (CD++, Algorithm 3) and enjoys an improved convergence that scales with $\sqrt{\bar{\kappa}_{k}}$ instead of $\bar{\kappa}_{k}$ (see Theorem 5.1). We test CD++ on a collection of benchmark positive definite problems from the machine learning literature [VvRBT13, PVG+11], which are known to exhibit large outlying eigenvalues. These experiments confirm our theoretical findings, showing that Kaczmarz methods can be competitive in floating point operations with both CG and GMRES on a range of input matrices.
@@
Further computational considerations and limitations.
While we focus our computational analysis on floating point operations, this metric may not fully capture the true running time, particularly when we wish to exploit sparsity in the input or parallelization in the hardware. Krylov subspace methods benefit from relying primarily on full matrix-vector product operations with ${\mathbf{A}}$ and can attain much faster performance, particularly for sparse and structured matrices. Below, we discuss how these considerations may affect the performance of Kaczmarz++ (and CD++).
  1. 1.
    Parallelization: While Kaczmarz methods often operate on single rows of the matrix at a time, Kaczmarz++ is specifically optimized for relying on large blocks of $\tilde{O}(k)$ rows. Since its dominant operation is typically an $\tilde{O}(k)\times n$ matrix-vector product, choosing a large $k$ not only improves the conditioning via $\bar{\kappa}_{k}$ but also helps the method take advantage of hardware parallelization.
  2. 2.
    Sparsity: A limitation of our theoretical analysis is that it only applies to solving dense linear systems. This is because randomized Hadamard preprocessing, which ensures an incoherence property that is needed for our theory, does not preserve the sparsity of the input. In Section 6, we show empirically that in some cases our methods still perform well without this step, since the input matrix ${\mathbf{A}}$ may already satisfy the needed incoherence property to begin with. Under these conditions, Kaczmarz++ can be implemented to run in time $\tilde{O}(nk^{2}+\mathrm{nnz}({\mathbf{A}})\bar{\kappa}_{k}\log 1/\epsilon)$, where $\mathrm{nnz}({\mathbf{A}})$ is the number of non-zeros in ${\mathbf{A}}$. In comparison, the cost of Krylov with orthogonalization to attain the same guarantee is $\tilde{O}(nT^{2}+\mathrm{nnz}({\mathbf{A}})T)$ for $T=O(k+\kappa_{k}\log 1/\epsilon)$. Nevertheless, future work (both theoretical and empirical) could assess the effectiveness of our algorithms on sparse problems.
  3. 3.
    Parameter tuning: The theoretically analyzed variant of Kaczmarz++ relies on hyper-parameters controlling momentum acceleration and block memoization. In the implemented variants of Kaczmarz++ and CD++, we use adaptive tuning to find these parameters during runtime. However, our algorithms still require correct selection of the block size (which is proportional to $k$) so that they can take full advantage of our guarantee (1.1). This stands in contrast to the Krylov guarantee (1.3) which holds for all $k$ simultaneously.
@@ 1.1 Overview of the Main Algorithm
Consider solving a consistent linear system with $m$ equations, ${\mathbf{A}}{\mathbf{x}}={\mathbf{b}}$, where ${\mathbf{A}}\in{\mathds{R}}^{m\times n}$ and ${\mathbf{b}}\in{\mathds{R}}^{m}$. The classical block Kaczmarz method constructs a sequence of iterates ${\mathbf{x}}_{0},{\mathbf{x}}_{1},...$ by repeatedly choosing a subset $S=S(t)\subseteq\{1,...,m\}=:[m]$ of those equations, and then projecting the current iterate ${\mathbf{x}}_{t}$ onto the subspace of solutions of those equations, i.e., the under-determined system ${\mathbf{A}}_{S}{\mathbf{x}}={\mathbf{b}}_{S}$. This leads to the block Kaczmarz update which can be stated as follows:
@@
We note that the Hadamard matrix, similarly to a Discrete Fourier Transform (DFT), is a scaled orthogonal matrix (specifically, ${\mathbf{H}}^{\scriptscriptstyle{\top}}{\mathbf{H}}=m{\mathbf{I}}_{m}$) that admits fast matrix-vector multiply (see Appendix E for a detailed discussion). In fact, these are the only two properties we use in our analysis, and one could replace Hadamard with a DFT or other fast transforms.
Since ${\mathbf{Q}}^{\scriptscriptstyle{\top}}{\mathbf{Q}}={\mathbf{I}}$, transforming the system ${\mathbf{A}}{\mathbf{x}}={\mathbf{b}}$ into ${\mathbf{Q}}{\mathbf{A}}{\mathbf{x}}={\mathbf{Q}}{\mathbf{b}}$ does not affect the solution of the system nor does it affect the singular value distribution of the input matrix. However, crucially, it ensures that all equations become roughly equally important (this is known as incoherence), which allows us to select a representative subset $S$ uniformly at random. Using the fast matrix-vector multiply, this transformation can be done using $mn\log m$ arithmetic operations. However, the resulting matrix ${\mathbf{Q}}{\mathbf{A}}$ does not retain the structural properties of ${\mathbf{A}}$ such as its sparsity pattern. Thus, while important for parts of our theoretical analysis, the RHT may be skipped when the input matrix is sparse and naturally exhibits an incoherence property [NT14].
@@
1:Input: ${\mathbf{A}}\in{\mathds{R}}^{m\times n}$, ${\mathbf{b}}\in{\mathds{R}}^{m}$, block size $s$, iterate ${\mathbf{x}}_{0}$, parameters $B$, ${\lambda}$, $\rho$, $\eta$;
@@ 1.2 Related work
Recently, there has been a number of works suggesting that variants of block Kaczmarz can exploit $k$ large outliers in the singular value distribution, by expressing its convergence in terms of $\bar{\kappa}_{k}$ or $\kappa_{k}$. [DR24] were the first to show this, however due to the large cost of Gaussian sketching and lack of acceleration, their method does not have a computational benefit over existing approaches. Then, [DY24] showed that a similar convergence guarantee can be obtained by block Kaczmarz with uniformly sampled blocks, after RHT preprocessing. Their algorithm converges in $\tilde{O}((n^{2}+nk^{2})\bar{\kappa}_{k}^{2}\log 1/\epsilon)$ operations. Most recently, [DLNR24] obtained $\tilde{O}((n^{2}+nk^{2})\kappa_{k}\log 1/\epsilon)$ operations by introducing momentum acceleration via a Nesterov-style scheme of [GHRS18] which is closely related to the momentum acceleration we use in Kaczmarz++. However, in order to provide a theoretical convergence analysis of their algorithm, [DLNR24] have to rely on a sketching-based block-construction strategy, which is much slower and less practical than uniform sampling. We resolve this theoretical limitation, and are able to use uniform block sampling in Kaczmarz++, by introducing regularized projections which facilitate our acceleration analysis, as discussed in Section 3.2.
@@ 1.3 Notation
[…]
@@ 2 Stable Convergence with Adaptive Acceleration
[…]
@@
Let $\rho\in[0,c\bar{\rho}]$ and $\eta\in[\frac{c}{\nu},\frac{1}{2\nu}]$ for some $c\in(0,1/2]$, and suppose that a sequence ${\mathbf{x}}_{t}$ is updated as in lines 6-10 of Algorithm 1, allowing the regularized projection step (line 8) to be computed inexactly, with $\tilde{\mathbf{w}}_{t}$ so that $\|\tilde{\mathbf{w}}_{t}-{\mathbf{w}}_{t}\|\leq\frac{\rho^{2}}{8\eta}\|{\mathbf{x}}_{t}-{\mathbf{x}}^{*}\|$. Then,
@@
If $\|\tilde{\mathbf{w}}_{t}-{\mathbf{w}}_{t}\|\leq\frac{\tilde{\mu}}{4}\|{\mathbf{x}}_{t}-{\mathbf{x}}^{*}\|$, then we have
@@ 3.1 Expectation of the Regularized Projection
[…]
@@
$\mu\Big{(}\bar{\mathbf{A}},{\mathcal{U}}(m,s),\lambda\Big{)}\geq\frac{\sigma_{\min}^{+}({\mathbf{A}})^{2}/2}{\sigma_{\min}^{+}({\mathbf{A}})^{2}+\bar{\lambda}}=\frac{1/2}{1+\frac{r-k}{k}\bar{\kappa}_{k}^{2}}\geq\frac{k}{2r\bar{\kappa}_{k}^{2}},$
@@
Remark 3.3.
The main part of our analysis that requires RHT is Lemma 3.6 below, from [DLNR24]. RHT preprocessing ensures that the input matrix ${\mathbf{A}}$ satisfies a deterministic incoherence condition which makes it amenable to uniform sub-sampling. For example, a sufficient (but not necessary) incoherence condition is that the matrix ${\mathbf{U}}$ of left singular vectors of ${\mathbf{A}}$ has all entries bounded as follows: $\max_{i,j}|u_{i,j}|=O(1/\sqrt{m})$. If this property holds for the input matrix ${\mathbf{A}}$, then Theorem 3.1 applies without RHT.
@@
Conditioned on the event $\mathcal{E}:=[S_{\mathrm{DPP}}\subseteq S]$ (which holds with probability $1-\delta^{\prime}$), we have ${\mathbf{A}}_{S_{\mathrm{DPP}}}^{\scriptscriptstyle{\top}}{\mathbf{A}}_{S_{\mathrm{DPP}}}\preceq{\mathbf{A}}_{S}^{\scriptscriptstyle{\top}}{\mathbf{A}}_{S}$. Combining this with (3.1) and Lemma 3.7, we have the following holds:
${\mathds{E}}[{\mathbf{P}}_{{\lambda},S}]=$ $~{\mathds{E}}[{\mathbf{P}}_{{\lambda},S}\mid\mathcal{E}]\Pr\{\mathcal{E}\}+{\mathds{E}}[{\mathbf{P}}_{{\lambda},S}\mid\neg\mathcal{E}]\Pr\{\neg\mathcal{E}\}$
$\succeq$ ${}\displaystyle~{\mathds{E}}[{\mathbf{P}}_{{\lambda},S}\mid\mathcal{E}]\Pr\{\mathcal{E}\}$
$\succeq$ ${}\displaystyle~\Pr\{\mathcal{E}\}\cdot{\mathbf{I}}-\frac{k\bar{\lambda}}{m}\cdot{\mathds{E}}\left[\left({\mathbf{A}}_{S}^{\scriptscriptstyle{\top}}{\mathbf{A}}_{S}+\frac{k\bar{\lambda}}{m}{\mathbf{I}}\right)^{-1}\biggl{|}\mathcal{E}\right]\cdot\Pr\{\mathcal{E}\}$
$\succeq$ ${}\displaystyle~\Pr\{\mathcal{E}\}\cdot{\mathbf{I}}-{\mathds{E}}\left[\left(\frac{m}{k\bar{\lambda}}{\mathbf{A}}_{S_{\mathrm{DPP}}}^{\scriptscriptstyle{\top}}{\mathbf{A}}_{S_{\mathrm{DPP}}}+{\mathbf{I}}\right)^{-1}\biggl{|}\mathcal{E}\right]\cdot\Pr\{\mathcal{E}\}$